OpenStudy (anonymous):

If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4). 17 −17 5 −5

OpenStudy (anonymous):

@Ashleyisakitty @brookie5150 @CoughCoughAchoo @dmezzullo @experimentX @Foreverisabell

OpenStudy (experimentx):

find f(x) and g(x) first

OpenStudy (anonymous):

I do not understand this. I have a lot of these so if you can explain it that would be a great help.

OpenStudy (anonymous):

how do you know what x or f or g are.

OpenStudy (experimentx):

f(x) = 2x − 2 put x=4 and find f(4)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

and g(4) is 11

OpenStudy (anonymous):

so i am finding (6-11)(4)

OpenStudy (anonymous):

I got -20 but that is not an option... did i do it wrong?

OpenStudy (experimentx):

yes ... find f(4) - g(4) ... that's correct

OpenStudy (anonymous):

but that is -20. My options are listed above so why does my answer not fit the options?

OpenStudy (experimentx):

check again ... you might have made error http://www.wolframalpha.com/input/?i=2x+%E2%88%92+2+-%28+3x+%E2%88%92+1%29+%2F.x-%3E4

OpenStudy (anonymous):

6-11=-5 -5x4=-20 where did i make a mistake????

OpenStudy (experimentx):

6-11=-5 <--- this is okay ... how did you get 20 ??

OpenStudy (anonymous):

-20 I then have to multiply it by four. that makes-20

OpenStudy (experimentx):

No no ... you don't have to multiply it by 4. (f-g)(x) = f(x) - g(x) <--- you misunderstood this.

OpenStudy (anonymous):

so my answer is -5.

OpenStudy (experimentx):

yes

OpenStudy (anonymous):

will you please help me with another? it is similar but has the parentheses in a different place and i am not sure how that effects it?

OpenStudy (experimentx):

sure ... post it.

OpenStudy (anonymous):

If f(x) = 3x + 1 and g(x) = 2x − 4, find f[g(2)]. 25 4 0 1

OpenStudy (experimentx):

find g(2) first.

OpenStudy (anonymous):

0

OpenStudy (anonymous):

is that all i have to do for this question because f is not getting multiplied by x.

OpenStudy (experimentx):

then find f(0)

OpenStudy (anonymous):

well that will be zero since anything times 0 is 0

OpenStudy (experimentx):

sorry ... f(0) = 3x0 + 1 = 1 so your answer is 1

OpenStudy (anonymous):

ok. thanks.

OpenStudy (experimentx):

yw ... what is f(x) = is not f times x ... okay, this is 'x' goes inside f

OpenStudy (anonymous):

ok. Now i understand.